OpenStudy (anonymous):

Sin[θ]= 0.7644 θ=???

OpenStudy (anonymous):

Perform algebra to remove the sin from the left side of the equation and leave only theta.

OpenStudy (anonymous):

The "opposite" of sine is inverse sine.

OpenStudy (anonymous):

\[ \theta = \arcsin(0.7644) \]This will give you the \[ -\pi/2 \leq \theta \leq \pi/2 \]solution. Then you want the \[ \pi/2 \leq \theta \leq 3\pi/2 \]solution.

OpenStudy (mrnood):

@1668168 This equation 'is' the answer - you cannot solve it any further by analytical means. You need to use a calculator or tables to find theta once you know sin(theta) on your calculator you will may have an 'inv' button - you need to press 'inv sin' OR you may have a sin^-1 button - which is the direct way to 'inverse sin' (make sure you have selected the calculator to work in degrees or radians - whichever is required by your answer...)